A certain man has a mass of 80 kg and a density of 955 kg/m3.


(a) Calculate his volume.

(b) Find the buoyant force air exerts on him.

(c) What is the ratio (in %) of the buoyant force to his weight?


Round your answer to the nearest 3 decimal places


1
Expert's answer
2021-04-19T17:07:39-0400

1) The volume is "V = \frac{m}{\rho} = \frac{80 kg}{955 \frac{kg}{m^3}} \approx 0.084 m^3".

2) The buoyant force is "F_b = \rho_a g V = 1.2 \frac{kg}{m^3} \cdot 9.81 \frac{m}{s^2} \cdot 0.084 m^3 \approx 0.99 N".

3) The ratio of buoyant force to weight is "\frac{F_b}{m g} = \frac{0.99 N}{80 kg \cdot 9.81 \frac{m}{s^2}} \approx 0.0013".